You are selling cups of coffee at a sporting event. Your sales records show that when the coffee is priced at $3 per cup, you sell 650 cups. When the coffee is prices at $2.50, you sell 930 cups.
Using that information, write a linear equation for the relationship between price (p) and quantity (q) sold. Your starting point is: q = mx + b, which is the slope intercept form of an equation, and using the two ordered pairs you find m and b. Make sure your answer here is an equation
Using the equation in #5 above, at what price of the coffee with you sell 800 cups? Round your answer to the nearest nickel, Use the formula Q=mP+b

Answer by stanbon(75887)

You are selling cups of coffee at a sporting event. Your sales records show that when the coffee is priced at $3 per cup, you sell 650 cups. When the coffee is prices at $2.50, you sell 930 cups.
----
(3,650);(2.5,*930)
-------------------------
Using that information, write a linear equation for the relationship between price (p) and quantity (q) sold. Your starting point is: q = mp + b, which is the slope intercept form of an equation, and using the two ordered pairs you find m and b. Make sure your answer here is an equation
---
m = (930-650)/(2.5-3) = 280/-0.5 = -560
Intercept ?
650 = -560*3 + b
b = 2330
Equation:
q(p) = -560p + 2330
-------------------------------
Using the equation in #5 above, at what price of the coffee with you sell 800 cups? Round your answer to the nearest nickel, Use the formula Q=mP+b
800 = -560p+2330
-560p = -1470
p = $2.63
=============
Cheers,
Stan H.

--------------------------------------------------------------------------------